With units of output 500, AVC $15, ATC $35, and MC $10, find total fixed cost, total variable cost, total cost and marginal cost.


1
Expert's answer
2021-03-26T08:08:29-0400
"FC=(ATC-AVC)\times Q=(35-15)\times500=10000"

"TC=ATC\times Q=35\times500=175000"

"C=MC\times Q=10\times 500=5000"